My age 5 years ago is half my age in 8 years

Answer by Cromlix(4381)

My age 5 years ago is half my age in 8 years
x - 5 = (x + 8)/2
Cross multiply
2(x - 5) = x + 8
2x - 10 = x + 8
Collect like terms;
2x - x = 8 + 10
x = 18
Your age is 18 years old.
Hope this helps.
:-)